Write all the factors of the following numbers: 
23 ????????????

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

23 = 1 × 23 23 = 23 × 1
Since 1 and 23 have occurred earlier
Hence, the factors of 23 are 1 and 23

Thanks